Can you please help me find an equation of the line containing the given pair of points?
(-7,0)and(0,5)
Thank you so much and God bless you!

Answer by RAY100(1637)

suggest simple sketch to clarify
slope = m=(y2-y1)/(x2-x1)= (5-0)/(0-(-7))= 5/7
in slope intercept form
y=mx+b
y=(5/7)x +b
subst (0,5)
5=(5/7) (0) +b
5=b
restating slope intercept form
y=mx+b
y=(5/7) x + 5 answer in slope intercept form
to convert to standard form (Ax+By=C)
multiple by 7
7y=5x +35
subst 5x both sides
-5x +7y=35
mult by (-1)
5x-7y=-35 answer in std form
check
(-7,0) 5(-7) -7(0) =-35 ok
(0,5) 5(0)-7(5)=-35 ok